OpenStudy (anonymous):

Two trees stand side-by-side. The shorter tree is 4 feet tall. Its shadow is 9 feet long. The taller tree is 18 feet tall. How long is its shadow? a) 4.5 b) 8 c) 23 d) 40.5

OpenStudy (cwrw238):

these will form two similar triangles so ratios of 2 heights = ratio of the 2 shadows 4 / 18 = 9 / x solve for x

OpenStudy (anonymous):

how?

OpenStudy (cwrw238):

cross multiply: 4x = 9*18 4x = 162 can you continue?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i got it thanks